OpenStudy (anonymous):

Secant = 5pi/6 find the exact values without a calculator or a unit circle. HELP!

OpenStudy (anonymous):

i'm assuming you mean\[sec(x) = \frac{ 5\pi }{ 6}\] and we want to solve for x, right?

OpenStudy (anonymous):

if its sec 5pi/6,then it is 5*180/6=5*30=150